# Enhancing inflationary model predictions via refined slow-roll dynamics

### v2

@@ 4 Numerical improvements
4.1 Implementation of accurate inflationary dynamics through numerical methods
@@
Figure 4: We plot the duration of reheating $N_{\rm re}$ and reheating temperature $T_{\rm re}$ given by Eqs. (3) and (3.9) as function of the scalar spectral index $n_{s}$ parametrically for both analytical and numerical solution, given by Eqs. (3.15), (2.16) respectively. The solid lines are for the numerical solution, and the dashed lines are for the analytically approximated solution. Please note that different colors represent dynamics corresponding to different effective equations of state parameter $w_{\rm re}$ as indicated in the figure. The blue-shaded region represents the $1\sigma$ constraint on the value of $n_{s}$ using ongoing observations [32, 33, 34, 35] with $n_{s}=0.9672\pm 0.0059$. The dark blue region shows the future projected bound on $n_{s}$ with a sensitivity of $10^{-3}$, assuming its central value remains unchanged. The temperature below the lighter red region is excluded due to the constraint from the electroweak scale, which is taken to be $100$ GeV.
Using leading-order slow-roll approximations and solving the slow-roll equations (viz. Eqs. (2.14)):
@@
along with the slow-roll solutions given in Eqs. (3.14). Here, $N_{\rm end}$ is the duration of the field rolling from the initial value $\phi_{i}$ to $\phi_{\rm end},$ i.e., the duration of inflation. These values are essential for the analysis in Eqs. (3) and (3.9) which has been shown in the previous section with Fig. 2.
To improve the analysis, in this section, we now consider, rather than the slow-roll, the full inflationary equations (viz. Eqs. (2.9) and (2.10)):
@@
Thus, there is a significant improvement not only in the numerical solution of the background variables but also in $H_{\rm end},$ $\rho_{\rm end},$ and $N_{\rm end}.$ These improvements directly impact the reheating analysis in two ways. The first is the direct influence of $\rho_{\rm end}$ in Eq. (3). The second, and more crucial, impact is the significant improvement in $\Delta N_{\rm end}=\Delta N_{k}\sim 1.5,$ which in turn enhances the accuracy of the background variables. For example, in the case of leading order approximations, $\{\epsilon_{1},\epsilon_{2}\}\simeq\{0.0003,0.040\}$ for $N_{k}=50,$ whereas, numerically we obtain these values as $\{\epsilon_{1},\epsilon_{2}\}\simeq\{0.00027,0.038\},$ and thus $\Delta n_{s}\simeq 2\times 10^{-3},$ as also discussed in Ref. [71]  [^2] Please note that, for a multi-parameter model, numerical evaluation for the entire range of the parameters can be challenging. In that case, one can rely upon a semi-analytical solution presented in Ref. [71], where analytically obtained solutions are nearly comparable to the numerical solutions..
Figure 5: We plot the inflationary observables, tensor-to-scalar ratio ($r$) as a function of scalar spectral index ($n_{s}$) for the Starobinsky model using numerical solution given by Eq. (2.16). This evolution is presented considering the bound on duration $N_{k}$ using the reheating regime. We consider different EoS parameters during reheating $w_{\rm re}$ and observe the constraint on the duration $N_{\rm k}$. Correspondingly, in the figure, blue line corresponds to the bound for $w_{\rm re}<1/3$, the gray line for $w_{\rm re}>1/3$ and the purple line for the future observational bound of $n_{\rm s}$. The blue-shaded region represents the $1\sigma$ constraint on the value of $n_{\rm s}$ using ongoing observations [32, 33, 34, 35] with $n_{\rm s}=0.9672\pm 0.0059$. The dark blue region shows the future projected bound on $n_{\rm s}$ with a sensitivity of $10^{-3}$, assuming its central value remains unchanged.
@@
Table 1: Starobinsky Inflation: The bounds on variables, $N_{k},~{}n_{s},~{}N_{\rm re},~{}\text{and}~{}T_{\rm re}$ for different values of $w_{\rm re}$ corresponding to analytically (slow-roll) approximated and the numerical solutions.
4.2 Implementation of higher-order slow-roll approximations
In the previous section, we estimated the reheating parameters and e-folding number using the numerical background solution with the leading order inflationary observables, i.e., Eqs. (2.16). However, as mentioned earlier, these expressions are obtained using leading-order slow-roll, and the full solutions of the amplitude of scalar power spectrum, scalar spectral index as well as the tensor-to-scalar ratio in terms of the slow-roll parameter are given as [73, 25]:
$A_{s}$ $=$ $1-2(C+1)\epsilon_{1}-C\epsilon_{2}+\left(2C^{2}+2C+\frac{\pi^{2}}{2}-f\right)\epsilon^{2}_{1}+\left(C^{2}-C+\frac{7\pi^{2}}{12}-g\right)\epsilon_{1}\epsilon_{2}$ (4.1)
@@ 4.2 Implementation of higher-order slow-roll approximations
$n_{s}$ $=$ $1-2\epsilon_{1}-\epsilon_{2}-2\epsilon^{2}_{1}-(3+2C)\epsilon_{1}\epsilon_{2}-C\epsilon_{2}\epsilon_{3},$ (4.2)
$r$ $=$ $16\epsilon_{1}(1+2C\frac{\epsilon_{2}}{2}),$ (4.3)
where $C=-2+\ln{2}+\gamma$, where $\gamma$ is the Euler constant, $f=5$ and $g=7$. In this section, instead of using Eqs. (2.16), along with the numerical solution, we consider the above expressions for the observable, repeat the reheating analysis, and compare our results with the numerical results obtained in the previous section, which has been illustrated in Figs. 6 and 7 as well as in Table 2.
Figure 6: We plot the duration of reheating $N_{\rm re}$ and reheating temperature $T_{\rm re}$ given by Eqs. (3) and (3.9) as function of the scalar spectral index $n_{s}$ parametrically comparing numerical solution with leading-order approximations and numerical solution with higher-order approximations given by Eq. (2.16), (4.2) respectively. The solid lines are for the higher-order approximations and the dashed lines are for the leading-order approximations. Please note that different colors represent dynamics corresponding to different effective equations of state parameter $w_{\rm re}$ as indicated in the figure. The blue-shaded region represents the $1\sigma$ constraint on the value of $n_{s}$ using ongoing observations[32, 33, 34, 35] with $n_{s}=0.9672\pm 0.0059$. The dark blue region shows the future projected bound on $n_{s}$ with a sensitivity of $10^{-3}$, assuming its central value remains unchanged. The temperature below the lighter red region is excluded due to the constraint from the electroweak scale, which is taken to be $100$ GeV.
@@
Figure 7: We plot the inflationary observables, tensor-to-scalar ratio ($r$) as a function of scalar spectral index ($n_{s}$) for the Starobinsky model using higher-order approximations given by Eq. (4.2), (4.3). This evolution is presented considering the bound on duration $N_{k}$ using the reheating regime. We consider different EoS parameters during reheating $w_{\rm re}$ and observe the constraint on the duration $N_{\rm k}$. Correspondingly, in the figure, blue line corresponds to the bound for $w_{\rm re}<1/3$, the gray line for $w_{\rm re}>1/3$ and the purple line for the future observational bound of $n_{\rm s}$. The blue-shaded region represents the $1\sigma$ constraint on the value of $n_{\rm s}$ using ongoing observations [32, 33, 34, 35] with $n_{\rm s}=0.9672\pm 0.0059$. The dark blue region shows the future projected bound on $n_{\rm s}$ with a sensitivity of $10^{-3}$, assuming its central value remains unchanged.
@@
Table 2: Starobinsky Inflation: The bounds on variables, $N_{k},~{}n_{s},~{}N_{\rm re},~{}\text{and}\ T_{\rm re}$ for different values of $w_{\rm re}$ corresponding to leading-order approximations and the higher-order approximations. Note that, in both cases, the background solutions are numerically obtained.
4.3 Implementation of the onset of reheating as the bottom of the potential
Till now, we were estimating $N_{\rm re}$ and $T_{\rm re}$, considering the onset of the reheating era after the end of inflation. However, in Ref. [62], it has been shown that, instead of considering minimal gravity theory, if one considers modified gravity theories, the onset of reheating as the end of inflation can a make a significant change in the analysis. This is because, $\epsilon_{1}$ is not a invariant under conformal transformation. Therefore, in a meaningful manner, one can choose the onset of reheating as the bottom of the potential, as this point is conformally invariant, and can be used even for general modified gravity theories. Therefore, in this section, we introduce corrections to the reheating analysis, and thus the observables, by considering that the reheating era commences after the epoch at which the potential touches its minima (bottom of the potential). In that case, Eq. (3) can be re-written as
$N_{\text{r}e}$ $=$ $\frac{4}{1-3w_{\text{re}}}\left(-N_{k}-N_{\text{eb}}-\frac{1}{4}\log\left(\frac{30}{\pi^{2}g_{\text{reh}}}\right)-\frac{1}{3}\log\left(\frac{11g_{\text{s,re}}}{43}\right)-\log\left(\frac{k}{a_{0}T_{0}}\right)\right.$ (4.4)
@@ 4.3 Implementation of the onset of reheating as the bottom of the potential
and the reheating temperature, i.e., Eq. (3.9) can also be evaluated as:
$T_{\text{re}}=\frac{a_{0}T_{0}}{k}\left(\frac{43}{11g_{\text{s,re}}}\right)e^{-N_{\text{k}}-N_{\text{eb}}-N_{\text{re}}}.$ (4.5)
@@
Figure 8: We plot the duration of reheating $N_{\rm re}$ and reheating temperature $T_{\rm re}$ given by Eqs. (3) and (3.9) as function of the scalar spectral index $n_{s}$ parametrically comparing analytically approximated solution and numerical solution with higher-order approximations, given by Eqs. (3.15), (4.2) respectively. The solid lines are for the onset of reheating as the bottom of the potential and the dashed lines are for the analytically approximated solution. Please note that different colors represent dynamics corresponding to different effective equations of state parameter $w_{\rm re}$ as indicated in the figure. The blue-shaded region represents the $1\sigma$ constraint on the value of $n_{s}$ using ongoing observations [32, 33, 34, 35] with $n_{s}=0.9672\pm 0.0059$. The dark blue region shows the future projected bound on $n_{s}$ with a sensitivity of $10^{-3}$, assuming its central value remains unchanged. The temperature below the lighter red region is excluded due to the constraint from the electroweak scale, which is taken to be $100$ GeV.
@@
Table 3: Starobinsky Inflation: The bounds on variables, $N_{k},~{}n_{s},~{}N_{\rm re},~{}\text{and}~{}T_{\rm re}$ for different values of $w_{\rm re}$ corresponding to analytical approximated parameters, and numerical solution with higher-order approximations.
5 Summary and conclusions
@@
    Numerical Solution: In this method, we solve the complete background Eq. (2.10) using the numerical method, which provides us with an accurate solution for the dynamics of the Universe in its early phase. Using this solution, we estimate the observables and compare the results with the solutions obtained using slow roll approximation.
@@
    Higher-order approximations: This method includes higher-order slow-roll corrections to the inflationary observables, Eqs. (4.1), (4.2) and (4.3), rather than Eqs. (2.16). Here also, we used complete numerical solution and estimated the observables using higher order slow-roll corrections and estimated the discrepancy in the observables with respect to the analytical solution obtained using slow-roll approximation.
@@
    Onset of the reheating as the bottom of the potential: Typically, the end of inflation is considered the onset of the reheating epoch. However, in this method, we show that if reheating begins at the bottom of the potential, the reheating parameters are modified, as given by Eq. (4.4) and Eq. (4.5). This modification can provide a new bound on the duration $N_{k}$, leading to a discrepancy in the scalar spectral index.

### v3

@@ IV Numerical improvements
IV.1 Implementation of accurate inflationary dynamics through numerical methods
@@
Figure 4: We plot the duration of reheating $N_{\rm re}$ and reheating temperature $T_{\rm re}$ given by Eqs. (III) and (25) as function of the scalar spectral index $n_{s}$ parametrically for both analytical and numerical solution, given by Eqs. (31) and (16), respectively. The solid lines are for the numerical solution, and the dashed lines are for the analytically approximated solution. Please note that different colors represent dynamics corresponding to different effective equations of state parameter $w_{\rm re}$ as indicated in the figure. The blue-shaded region represents the $1\sigma$ constraint on the value of $n_{s}$ using ongoing observations [32, 33, 34, 35] with $n_{s}=0.9672\pm 0.0059$. The dark blue region shows the future projected bound on $n_{s}$ with a sensitivity of $10^{-3}$, assuming its central value remains unchanged. The temperature below the lighter red region is excluded due to the constraint from the electroweak scale, which is taken to be 100100 GeV.
Using leading-order slow-roll approximations and solving the slow-roll equations (viz. Eqs. (14)):
@@
along with the slow-roll solutions given in Eqs. (30). Here, $N_{\rm end}$ is the duration of the field rolling from the initial value $\phi_{i}$ to $\phi_{\rm end},$ i.e., the duration of inflation. These values are essential for the analysis in Eqs. (III) and (25) which has been shown in the previous section with Fig. 2.
To improve the analysis, in this section, we now consider, rather than the slow-roll, the full inflationary equations (viz. Eqs. (9) and (10)):
@@
Thus, there is a significant improvement not only in the numerical solution of the background variables but also in $H_{\rm end},$ $\rho_{\rm end},$ and $N_{\rm end}.$ These improvements directly impact the reheating analysis in two ways. The first is the direct influence of $\rho_{\rm end}$ in Eq. (III). The second, and more crucial, impact is the significant improvement in $\Delta N_{\rm end}=\Delta N_{k}\sim 1.5,$ which in turn enhances the accuracy of the background variables. For example, in the case of leading order approximations, $\{\epsilon_{1},\epsilon_{2}\}\simeq\{0.0003,0.040\}$ for $N_{k}=50,$ whereas, numerically we obtain these values as $\{\epsilon_{1},\epsilon_{2}\}\simeq\{0.00027,0.038\},$ and thus $\Delta n_{s}\simeq 2\times 10^{-3},$ as also discussed in Ref. [72]  [^2] Please note that, for a multi-parameter model, numerical evaluation for the entire range of the parameters can be challenging. In that case, one can rely upon a semi-analytical solution presented in Ref. [72], where analytically obtained solutions are nearly comparable to the numerical solutions..
Figure 5: We plot the inflationary observables, tensor-to-scalar ratio ($r$) as a function of scalar spectral index ($n_{s}$) for the Starobinsky model using numerical solution given by Eq. (16). This evolution is presented considering the bound on duration $N_{k}$ using the reheating regime. We consider different EoS parameters during reheating $w_{\rm re}$ and observe the constraint on the duration $N_{\rm k}$. Correspondingly, in the figure, blue line corresponds to the bound for $w_{\rm re}<1/3$, the gray line for $w_{\rm re}>1/3$ and the purple line for the future observational bound of $n_{\rm s}$. The blue-shaded region represents the $1\sigma$ constraint on the value of $n_{\rm s}$ using ongoing observations [32, 33, 34, 35] with $n_{\rm s}=0.9672\pm 0.0059$. The dark blue region shows the future projected bound on $n_{\rm s}$ with a sensitivity of $10^{-3}$, assuming its central value remains unchanged.
@@
Table 1: Starobinsky Inflation: The bounds on variables $N_{k},~n_{s},~N_{\rm re},~\text{and}~T_{\rm re}$ for different values of $w_{\rm re}$ corresponding to analytically (slow-roll) approximated and the numerical solutions are shown.
IV.2 Implementation of higher-order slow-roll approximations
In the previous section, we estimated the reheating parameters and e-folding number using the numerical background solution with the leading order inflationary observables, i.e., Eqs. (16). However, as mentioned earlier, these expressions are obtained using leading-order slow-roll, and the full solutions of the amplitude of scalar power spectrum, scalar spectral index as well as the tensor-to-scalar ratio in terms of the slow-roll parameter are given as [74, 25]:
$A_{s}$ $=$ $1-2(C+1)\epsilon_{1}-C\epsilon_{2}+\left(2C^{2}+2C+\frac{\pi^{2}}{2}-f\right)\epsilon^{2}_{1}+\left(C^{2}-C+\frac{7\pi^{2}}{12}-g\right)\epsilon_{1}\epsilon_{2}$ (38)
@@ IV.2 Implementation of higher-order slow-roll approximations
$n_{s}$ $=$ $1-2\epsilon_{1}-\epsilon_{2}-2\epsilon^{2}_{1}-(3+2C)\epsilon_{1}\epsilon_{2}-C\epsilon_{2}\epsilon_{3},$ (39)
$r$ $=$ $16\epsilon_{1}(1+2C\frac{\epsilon_{2}}{2}),$ (40)
where $C=-2+\ln{2}+\gamma$, where $\gamma$ is the Euler constant, $f=5$ and $g=7$. In this section, instead of using Eqs. (16), along with the numerical solution, we consider the above expressions for the observable, repeat the reheating analysis, and compare our results with the numerical results obtained in the previous section, which has been illustrated in Figs. 6 and 7 as well as in Table 2.
Figure 6: We plot the duration of reheating $N_{\rm re}$ and reheating temperature $T_{\rm re}$ given by Eqs. (III) and (25) as function of the scalar spectral index $n_{s}$ parametrically comparing numerical solution with leading-order approximations and numerical solution with higher-order approximations given by Eq. (16), (39) respectively. The solid lines are for the higher-order approximations and the dashed lines are for the leading-order approximations. Please note that different colors represent dynamics corresponding to different effective equations of state parameter $w_{\rm re}$ as indicated in the figure. The blue-shaded region represents the $1\sigma$ constraint on the value of $n_{s}$ using ongoing observations[32, 33, 34, 35] with $n_{s}=0.9672\pm 0.0059$. The dark blue region shows the future projected bound on $n_{s}$ with a sensitivity of $10^{-3}$, assuming its central value remains unchanged. The temperature below the lighter red region is excluded due to the constraint from the electroweak scale, which is taken to be 100100 GeV.
@@
Figure 7: We plot the inflationary observables, tensor-to-scalar ratio ($r$) as a function of scalar spectral index ($n_{s}$) for the Starobinsky model using higher-order approximations given by Eq. (39) and (40). This evolution is presented considering the bound on duration $N_{k}$ using the reheating regime. We consider different EoS parameters during reheating $w_{\rm re}$ and observe the constraint on the duration $N_{\rm k}$. Correspondingly, in the figure, blue line corresponds to the bound for $w_{\rm re}<1/3$, the gray line for $w_{\rm re}>1/3$ and the purple line for the future observational bound of $n_{\rm s}$. The blue-shaded region represents the $1\sigma$ constraint on the value of $n_{\rm s}$ using ongoing observations [32, 33, 34, 35] with $n_{\rm s}=0.9672\pm 0.0059$. The dark blue region shows the future projected bound on $n_{\rm s}$ with a sensitivity of $10^{-3}$, assuming its central value remains unchanged.
@@
Table 2: Starobinsky Inflation: The bounds on variables $N_{k},~n_{s},~N_{\rm re},~\text{and}\ T_{\rm re}$ for different values of $w_{\rm re}$ corresponding to leading-order approximations and the higher-order approximations are shown. Note that, in both cases, the background solutions are numerically obtained.
IV.3 Implementation of the onset of reheating as the bottom of the potential
Till now, we were estimating $N_{\rm re}$ and $T_{\rm re}$, considering the onset of the reheating era after the end of inflation. However, in Ref. [63], it has been shown that, instead of considering minimal gravity theory, if one considers modified gravity theories, the onset of reheating as the end of inflation can a make a significant change in the analysis. This is because, $\epsilon_{1}$ is not a invariant under conformal transformation. Therefore, in a meaningful manner, one can choose the onset of reheating as the bottom of the potential, as this point is conformally invariant, and can be used even for general modified gravity theories. Therefore, in this section, we introduce corrections to the reheating analysis, and thus the observables, by considering that the reheating era commences after the epoch at which the potential touches its minima (bottom of the potential). In that case, Eq. (III) can be re-written as
$N_{\text{r}e}$ $=$ $\frac{4}{1-3w_{\text{re}}}\left(-N_{k}-N_{\text{eb}}-\frac{1}{4}\log\left(\frac{30}{\pi^{2}g_{\text{reh}}}\right)-\frac{1}{3}\log\left(\frac{11g_{\text{s,re}}}{43}\right)-\log\left(\frac{k}{a_{0}T_{0}}\right)\right$. (41)
@@ IV.3 Implementation of the onset of reheating as the bottom of the potential
and the reheating temperature, i.e., Eq. (25) can also be evaluated as:
$T_{\text{re}}=\frac{a_{0}T_{0}}{k}\left(\frac{43}{11g_{\text{s,re}}}\right)e^{-N_{\text{k}}-N_{\text{eb}}-N_{\text{re}}}.$ (42)
@@
Figure 8: We plot the duration of reheating $N_{\rm re}$ and reheating temperature $T_{\rm re}$ given by Eqs. (III) and (25) as function of the scalar spectral index $n_{s}$ parametrically comparing analytically approximated solution and numerical solution with higher-order approximations, given by Eqs. (31), (39) respectively. The solid lines are for the onset of reheating as the bottom of the potential and the dashed lines are for the analytically approximated solution. Please note that different colors represent dynamics corresponding to different effective equations of state parameter $w_{\rm re}$ as indicated in the figure. The blue-shaded region represents the $1\sigma$ constraint on the value of $n_{s}$ using ongoing observations [32, 33, 34, 35] with $n_{s}=0.9672\pm 0.0059$. The dark blue region shows the future projected bound on $n_{s}$ with a sensitivity of $10^{-3}$, assuming its central value remains unchanged. The temperature below the lighter red region is excluded due to the constraint from the electroweak scale, which is taken to be 100100 GeV.
@@
Table 3: Starobinsky Inflation: The bounds on variables, $N_{k},~n_{s},~N_{\rm re},~\text{and}~T_{\rm re}$ for different values of $w_{\rm re}$ corresponding to analytical approximated parameters, and numerical solution with higher-order approximations are shown.
V Summary and conclusions
@@
    Numerical Solution: In this method, we solve the complete background Eq. (10) using the numerical method, which provides us with an accurate solution for the dynamics of the universe in its early phase. Using this solution, we estimate the observables and compare the results with the solutions obtained using slow roll approximation.
@@
    Higher-order approximations: This method includes higher-order slow-roll corrections to the inflationary observables, Eqs. (38), (39) and (40), rather than Eqs. (16). Here also, we used complete numerical solution and estimated the observables using higher order slow-roll corrections and estimated the discrepancy in the observables with respect to the analytical solution obtained using slow-roll approximation.
@@
    Onset of the reheating as the bottom of the potential: Typically, the end of inflation is considered the onset of the reheating epoch. However, in this method, we show that if reheating begins at the bottom of the potential, the reheating parameters are modified, as given by Eq. (41) and Eq. (42). This modification can provide a new bound on the duration $N_{k}$, leading to a discrepancy in the scalar spectral index.
@@
Acknowledgements
DN is supported by the DST, Government of India through the DST-INSPIRE Faculty fellowship (04/2020/002142). MK is supported by a DST-INSPIRE Fellowship under the reference number: IF170808, DST, Government of India. DN, SY and MK are also very thankful to the Department of Physics and Astrophysics, University of Delhi. MK, SY and DN also acknowledge facilities provided by the IUCAA Centre for Astronomy Research and Development (ICARD), University of Delhi.